how high is a tree that casts a 33 ft shadow at the same time a 5 ft pole casts a shadow which is 12 ft long?
Answer by lwsshak3(11628)

how high is a tree that casts a 33 ft shadow at the same time a 5 ft pole casts a shadow which is 12 ft long?
let x=height of the tree
x/33=5/12
x=5*33/12
x=13.75
height of the tree=13.75 ft